# On the Dirichlet problem for the degenerate $k$-Hessian equation

### v1

@@ Abstract.
This paper investigates the existence of a global $C^{1,1}$ solution to the Dirichlet problem for the $k$-Hessian equation with a nonnegative right-hand side $f$, focusing on the required conditions for $f$. A well-known counterexample demonstrates the sharpness of the condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in a domain $\Omega_{0}\Supset\Omega$. We establish the existence for all $2\leq k\leq n-1$ under the condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$, and under an a priori assumption that $\Delta u\geq 1$. Surprisingly, higher smoothness of $f$ permits a better exponent to be achieved. The condition $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$ is also sharp, as demonstrated by the same counterexample. For the Monge-Ampère equation (the case $k=n$), we establish the existence under the optimal condition $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$. For any $5\leq k\leq n-1$, we prove the existence under the condition $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$, and under the a priori assumption $\Delta u\geq 1$. Moreover, we obtain the existence for all $2\leq k\leq n-1$ under the condition $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$ with $f\geq 0$ in $\Omega_{0}$.
@@ 1. Introduction
In this paper, we consider the Dirichlet problem for the degenerate $k$-Hessian equation:
@@
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ ($n\geq 2$), $2\leq k\leq n$, and $D^{2}u$ denotes the Hessian matrix of the function $u$. For an $n\times n$ real symmetric matrix $A$, we denote by $\sigma_{k}[A]$ the $k$-th elementary symmetric polynomial of its eigenvalues $\lambda(A)=(\lambda_{1},\dots,\lambda_{n})$. That is,
$\sigma_{k}[A]=\sigma_{k}(\lambda(A)):=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}},\quad\text{for}\ k=1,2,\dots,n.$
Equivalently, $\sigma_{k}[A]$ is the sum of all the $k\times k$ principal minors of $A$. By convention, we set $\sigma_{0}(\lambda)\equiv 1$.
@@
$\lambda\big(D^{2}u(x)\big)\in\Gamma_{k}\quad\text{for all}\ x\in\Omega,$
where $\Gamma_{k}$ denotes the Gårding cone defined by
$\Gamma_{k}:=\left\{\lambda\in\mathbb{R}^{n}:\ \sigma_{j}(\lambda)>0\ \text{for}\ j=1,2,\dots,k\right\}.$
They established that for any $\lambda\in\Gamma_{k}$ and each $1\leq i\leq n$,
$\frac{\partial\sigma_{k}}{\partial\lambda_{i}}>0;$
that $\sigma_{k}^{1/k}(\lambda)$ is a concave function in $\Gamma_{k}$; and that
$\text{the operator}\ \,\sigma_{k}^{1/k}[A]\,\ \text{is elliptic and concave in}\ \ \left\{A\in\mathscr{S}^{n}:\ \lambda(A)\in\Gamma_{k}\right\},$
where $\mathscr{S}^{n}$ denotes the space of $n\times n$ real symmetric matrices. A hypersurface in $\mathbb{R}^{n}$ is called $k$-convex (for $k\in\{1,2,\dots,n-1\}$) if its principal curvatures $\kappa=(\kappa_{1},\kappa_{2},\dots,\kappa_{n-1})$ satisfy $\sigma_{j}(\kappa)\geq 0$ for all $j\leq k$ everywhere. It is strictly $k$-convex if $\sigma_{j}(\kappa)\geq\delta>0$ for some constant $\delta$ uniformly on the hypersurface for all $j\leq k$. As shown in [2], a necessary condition for the existence of a $C^{2}(\overline{\Omega})$ solution to (1.1) is that the boundary $\partial\Omega$ be strictly (k1)(k-1)-convex when viewed as a hypersurface in $\mathbb{R}^{n}$.
@@
Wang [32] proved that if $\beta<2(n-1)(\alpha-1)$, then (1.2) admits no $C^{1,1}$ solution. Taking $\beta=2(n-1)(\alpha-1)-1$, we have $\bar{f}\geq 0$ in $B_{2}$, and
@@
Ivochkina-Trudinger-Wang [16] pointed out that a modification of the counterexample (1.2) also applies to the $k$-Hessian equations for $2\leq k\leq n-1$, although no proof was provided. A detailed proof of this fact was given by Dinew-Pliś-Zhang [5]; namely, the problem
@@
are two best possible conditions.
For the Monge-Ampère equation, Guan [9] reduced the estimate of second-order derivatives to the boundary under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, and obtained the existence of a convex solution in $C^{1,1}(\overline{\Omega})$ under the homogeneous boundary condition ($\varphi=0$). Subsequently, by introducing a completely new method, Guan-Trudinger-Wang [12] established a global $C^{2}$ estimate for the Monge-Ampère equation under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, which led to the existence of a convex solution in $C^{1,1}(\overline{\Omega})$. For the case $2\leq k\leq n-1$ of equation (1.1), Ivochkina-Trudinger-Wang [16] proposed the following open problem:
[…]
Dong [6] solved the problem (1.3) under the homogeneous boundary condition. For general boundary conditions, Jiao-Wang [17] derived a global $C^{2}$ estimate for convex solutions to problem (1.3). However, since smooth approximations of convex viscosity solutions do not necessarily preserve convexity, they [17] were unable to prove the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for equation (1.1).
In this paper, we investigate for equation (1.1) both problem (1.3) and the following problem: for any $k\in\{2,3,\dots,n\}$,
[…]
where the exponent p=2(k1)/3p=2(k-1)/3 is best possible. We solve problem (1.4) for the Monge-Ampère equation under the optimal condition p=2(n1)/3p=2(n-1)/3, and for general $k$-Hessian equations ($2\leq k\leq n-1$) with $p=2k/3$. Furthermore, we address problem (1.3) and problem (1.4) ($5\leq k\leq n-1$) under their respective best possible conditions, and under the following additional a priori assumption: there exist small positive constants $\delta_{0}$ and $r_{0}$ such that
$\Delta u\geq\delta_{0}\quad\text{in}\ \Omega\cap B_{r_{0}}(y),$ (1.5)
for all $y\in\partial\Omega$ at which $\Delta u(y)\gg 1$.
In order to find a global $C^{1,1}$ $k$-admissible solution in degenerate problems, the usual strategy is to establish global $C^{2}$ a priori estimates for the non-degenerate case that are independent of the positive lower bound of $f$. Then, based on the existence theory of Caffarelli-Nirenberg-Spruck [2] for the non-degenerate case, the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for the degenerate case is obtained by an approximation argument. There is an extensive body of work on this problem. In addition to the references cited, we refer the reader to [3, 15, 18, 19, 28, 29, 30, 11, 8, 13, 31, 25, 4]. This is, however, a far from complete list of important contributions.
Below we state the main results. First, we establish global $C^{1}$ estimate and boundary estimates for the second-order derivatives in the tangential and tangential-normal directions.
@@
$\|u\|_{C^{1}(\overline{\Omega})}+\sup_{\partial\Omega}\left(|u_{\nu\eta}|+|u_{\eta\eta}|\right)\leq C,$
where $\nu$ denotes the unit interior normal vector, $\eta$ is any unit tangential vector; and $C$ depends on $\Omega$, $\|\varphi\|_{C^{3}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k-2)}\|_{C^{1,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Second, we reduce the estimate of second-order derivatives to the boundary.
Theorem 1.2.
Under the homogeneous boundary condition, Dong [6] established the required second-order normal derivative estimate on the boundary. Combining this result with Theorems 1.1 and 1.2 yields the global $C^{2}$ estimate.
Theorem 1.3.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi=0$, $\inf_{\Omega_{0}}f>0$, and $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
The existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.4.
Results corresponding to Theorems 1.1, 1.2, 1.3 and Corollary 1.4 were established by Jiao-Wang [17] under the weaker condition $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega})$ with $f\geq 0$ in $\Omega$. In this paper, we focus on problems (1.3) and (1.4) with general boundary conditions. The primary challenge, therefore, is to estimate the second-order derivatives in the normal direction on the boundary. For general $k$-Hessian equations, we employ the framework of Ivochkina-Trudinger-Wang [16] to establish the boundary second-order normal derivative estimates.
Ivochkina-Trudinger-Wang [16] provided a PDE-based proof of Krylov’s results [21, 22, 23, 24], and roughly speaking, their ideas are similar. These results and ideas were clarified very well in [16], and we quote them below.
@@
Third, we establish the weakly interior estimate for degenerate $k$-Hessian equations.
@@
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. For any constant $\varepsilon>0$, there exists a positive constant $C_{\varepsilon}$ such that
@@
Fourth, we obtain the boundary estimate in terms of the interior one for degenerate $k$-Hessian equations under the condition (1.5). Then the global $C^{2}$ estimate is established.
@@
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Let the condition (1.5) hold. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where the constant $C$ depends on $\delta_{0}$, $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and either $\|f^{1/(k-1)}\|_{C^{1,1}(\overline{\Omega_{0}})}$ or $\|f^{3/(2k-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
@@
The condition (1.5) is naturally satisfied if $f$ has a positive lower bound near $\partial\Omega$.
The existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.8.
Let $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, and $f\geq 0$ in $\Omega_{0}$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Let the condition (1.5) hold. Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Fifth, if $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$, then the condition (1.5) can be omitted. As pointed out by Guan-Trudinger-Wang [12], for applications to problems in differential geometry, it is desirable to impose no restriction on the nonnegative function $f$ apart from smoothness. Guan [9] established the existence of a convex solution in $C^{1,1}(\overline{\Omega})$ for the Monge-Ampère equation with the homogeneous boundary condition when $n=3$ and $0\leq f\in C^{3,1}(\mathbb{R}^{3})$. Compared to the case when $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$, it is surprising that the exponent of $f$ can be improved provided its smoothness is higher, for $k\geq 4$.
Theorem 1.9.
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$ and $\inf_{\Omega_{0}}f>0$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where the constant $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2k)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
The existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.10.
Let $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}})$ and $f\geq 0$ in $\Omega_{0}$, where $\Omega\Subset\Omega_{0}$. Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Remark 1.11.
There is no inclusion relationship between the conditions $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$ and $f^{1/(k-q)}\in C^{2,1}(\overline{\Omega_{0}})$ for any $0<q<k$. This is demonstrated by two counterexamples. One counterexample is $f^{1/(k-q)}:=|x|^{2}\in C^{\infty}(\overline{\Omega_{0}})$, where the origin $0\in\partial\Omega$. It is clear that $f^{1/k}=|x|^{2(k-q)/k}\notin C^{1,1}(\overline{\Omega_{0}})$. The other counterexample is
@@
in $\overline{\Omega_{0}}$, where the origin 0 and $x_{0}$ are two distinct points on $\partial\Omega$, and
$g(x):=\begin{cases}|x|^{4}\left(1+\sin\frac{1}{|x|}\right),\quad x\neq 0,\\ 0,\quad x=0.\end{cases}$
One can verify that $g\in W^{2,\infty}$ which implies $g\in C^{1,1}(\overline{\Omega_{0}})$, and that $D^{2}g$ is not continuous at the origin. Consequently, $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$, but $D^{2}(f^{1/(k-q)})$ fails to be continuous at the origin.
For the Monge-Ampère equation, using the convexity of the admissible solutions and of the domain, as well as the affine invariance of the equation, Guan-Trudinger-Wang [12] introduced a completely new method to estimate the second-order normal derivatives on the boundary. Recently, Jiao-Wang [17] extended this method to convex solutions of the $k$-Hessian equation in uniformly convex domains. Sixth, applying the method of Guan-Trudinger-Wang [12], we solve problem (1.4) for the Monge-Ampère equation under the optimal condition.
Theorem 1.12.
Let $u\in C^{3,1}(\overline{\Omega})$ be a convex solution of equation (1.1) with $k=n$, $\partial\Omega\in C^{3,1}$ be uniformly convex, $\varphi\in C^{3,1}(\partial\Omega)$, $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $\inf_{\Omega_{0}}f>0$, where $\Omega\Subset\Omega_{0}$. Then
$\|u\|_{C^{2}(\overline{\Omega})}\leq C,$
where the constant $C$ depends on $\Omega$, $\|\varphi\|_{C^{3,1}(\partial\Omega)}$, $\operatorname{dist}(\Omega,\partial\Omega_{0})$, and $\|f^{3/(2n-2)}\|_{C^{2,1}(\overline{\Omega_{0}})}$, but is independent of $\inf_{\Omega_{0}}f$.
Remark 1.13.
The condition in $\Omega_{0}$ of Theorem 1.12 (rather than in $\Omega$) is only used to reduce the estimate of second-order derivatives to the boundary, as established in Theorem 1.2.
The existence of a convex solution in $C^{1,1}(\overline{\Omega})$ is established by an approximation argument using solutions to non-degenerate problems.
Corollary 1.14.
The rest of the paper is organized as follows. Section 2 presents preliminary results. Section 3 establishes the global $C^{1}$ estimate and the boundary second-order tangential and tangential-normal estimates (Theorem 1.1). Section 4 reduces the estimate of second-order derivatives to the boundary (Theorem 1.2). Section 5 is devoted to the weakly interior estimate, namely Theorem 1.5. Section 6 develops the boundary estimate in terms of the interior one to finish Theorems 1.6 and 1.9. Section 7 solves problem (1.4) for the Monge-Ampère equation under the optimal condition, namely Theorem 1.12. The Appendix (Section 8) contains auxiliary lemmas.
In this paper, the dependence of constants on dimension $n$ and the number $k$ in (1.1) is omitted, and a constant depending only on $n$ and $k$ is called universal. This section ends by listing some notations.
@@
10. $D^{2}_{x^{\prime}x^{\prime}}$ denotes the $(n-1)\times(n-1)$ matrix with elements $\partial_{ij}$ where $1\leq i,\ j\leq n-1$.

### v2

@@ Abstract.
This paper investigates the existence of a global $C^{1,1}$ solution to the Dirichlet problem for the $k$-Hessian equation with a nonnegative right-hand side $f$, focusing on the required conditions for $f$. The conditions $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ and $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$, together with $f\geq 0$ in a domain $\Omega_{0}\Supset\Omega$, are optimal, as demonstrated by classical counterexamples. For the Monge-Ampère equation ($k=n$), we establish the existence under the optimal condition $f^{3/(2n-2)}\in C^{2,1}(\overline{\Omega_{0}})$ together with $f\geq 0$ in $\Omega_{0}$. For the general $k$-Hessian equation, we establish the existence under the condition $f\geq 0$ in $\Omega_{0}$ together with one of the following three conditions:
$(i)\quad f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}}),\ \ \inf_{\Omega}\Delta u\geq 1,\ \ 2\leq k\leq n-1;$
$(ii)\quad f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}}),\ \ \inf_{\Omega}\Delta u\geq 1,\ \ 5\leq k\leq n-1;$
$(iii)\quad f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}}),\ \ 2\leq k\leq n-1.$
@@ 1. Introduction
In this paper, we consider the Dirichlet problem for the degenerate $k$-Hessian equation
@@
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$, $2\leq k\leq n$, and $D^{2}u$ denotes the Hessian matrix of the function $u$. For an $n\times n$ real symmetric matrix $A$, we denote by $\sigma_{k}[A]$ the $k$-th elementary symmetric polynomial of the eigenvalues $\lambda(A)=(\lambda_{1},\dots,\lambda_{n})$; that is,
$\sigma_{k}[A]=\sigma_{k}(\lambda(A)):=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}$
for $k=1,2,\dots,n$. Equivalently, $\sigma_{k}[A]$ is the sum of all $k\times k$ principal minors of $A$.
@@
$\lambda\big(D^{2}u(x)\big)\in\Gamma_{k},\quad\forall x\in\Omega,$
where the Gårding cone is defined by
$\Gamma_{k}:=\left\{\lambda\in\mathbb{R}^{n}:\ \sigma_{j}(\lambda)>0\ \text{for all}\ 1\leq j\leq k\right\}.$
They proved that
$\frac{\partial\sigma_{k}}{\partial\lambda_{i}}>0\quad\text{in}\ \Gamma_{k},\quad\forall 1\leq i\leq n;$
that $\sigma_{k}^{1/k}(\lambda)$ is concave in $\Gamma_{k}$; and that the operator $\sigma_{k}^{1/k}[A]$ is elliptic and concave for any real symmetric matrix $A$ with $\lambda(A)\in\Gamma_{k}$. A hypersurface in $\mathbb{R}^{n}$ is called $k$-convex (for $k\in\{1,2,\dots,n-1\}$) if its principal curvatures $\kappa=(\kappa_{1},\kappa_{2},\dots,\kappa_{n-1})$ satisfy $\sigma_{j}(\kappa)\geq 0$ for all $j\leq k$ everywhere. It is strictly $k$-convex if there exists a constant $\delta>0$ such that $\sigma_{j}(\kappa)\geq\delta$ uniformly on the hypersurface for all $j\leq k$. As shown in [2], a necessary condition for the existence of a $C^{2}(\overline{\Omega})$ solution to (1.1) is that the boundary $\partial\Omega$ be strictly (k1)(k-1)-convex when regarded as a hypersurface in $\mathbb{R}^{n}$.
@@
Wang [32] proved that if $\beta<2(n-1)(\alpha-1)$, then (1.2) admits no $C^{1,1}$ solution. Taking $\beta=2(n-1)(\alpha-1)-1$, one has $\bar{f}\geq 0$ in $B_{2}$, and
@@
Ivochkina-Trudinger-Wang [16] pointed out that a modification of the counterexample (1.2) also applies to the $k$-Hessian equation for $2\leq k\leq n-1$, although no proof was provided. A detailed proof of this fact was given by Dinew-Pliś-Zhang [5]; namely, the problem
@@
are optimal.
For the Monge-Ampère equation, Guan [9] reduced the estimate of second-order derivatives to the boundary under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, and obtained the existence of a convex solution in $C^{1,1}(\overline{\Omega})$ under the homogeneous boundary condition ($\varphi=0$). Subsequently, by introducing a completely new method, Guan-Trudinger-Wang [12] established a global $C^{2}$ estimate for the Monge-Ampère equation under the optimal condition $f^{1/(n-1)}\in C^{1,1}(\overline{\Omega})$, which in turn yielded the existence of a convex solution in $C^{1,1}(\overline{\Omega})$. For the case $2\leq k\leq n-1$ of equation (1.1), Ivochkina-Trudinger-Wang [16] proposed the following open problem:
$\text{existence of a global $C^{1,1}$ solution under $f^{1/(k-1)}\in C^{1,1}$}.$ (1.3)
Dong [6] solved problem (1.3) under the homogeneous boundary condition. For general boundary conditions, Jiao-Wang [17] derived a global $C^{2}$ estimate for convex solutions to problem (1.3). However, since smooth approximations of convex viscosity solutions do not necessarily preserve convexity, they [17] were unable to establish the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for equation (1.1). There is an extensive literature on degenerate $k$-Hessian equations; see, for instance, [3, 15, 18, 19, 28, 29, 30, 11, 8, 13, 31, 25, 4], although this list is far from exhaustive.
A common strategy is to derive global $C^{2}$ a priori estimate for the non-degenerate case that are independent of the positive lower bound of $f$. Then the existence of a $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for the degenerate case is established by an approximation argument using solutions to non-degenerate problems. To establish global $C^{2}$ a priori estimate independent of the positive lower bound of $f$, the primary challenge is to derive the second-order normal derivative estimate on the boundary.
For the Monge-Ampère equation, using the convexity of the admissible solutions and of the domain, as well as the affine invariance of the equation, Guan-Trudinger-Wang [12] introduced a method to estimate the second-order normal derivatives on the boundary. Recently, Jiao-Wang [17] extended this method to convex solutions of the $k$-Hessian equation in uniformly convex domains. Applying the method of Guan-Trudinger-Wang [12], we establish the existence of a global $C^{1,1}$ solution for the Monge-Ampère equation under the optimal condition.
@@
For general $k$-Hessian equations, we employ the method of Ivochkina-Trudinger-Wang [16], which is completely different from that of [12], to establish the boundary second-order normal derivative estimate. Ivochkina-Trudinger-Wang [16] provided a PDE-based proof of Krylov’s results [21, 22, 23, 24], and, roughly speaking, their ideas are similar. These results and ideas were clarified very well in [16], and we quote them below.
@@
Firstly, we establish the weakly interior estimate for degenerate $k$-Hessian equations.
@@
Let $u\in C^{3,1}(\overline{\Omega})$ be a $k$-admissible solution of equation (1.1), $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, $\inf_{\Omega_{0}}f>0$ where $\Omega\Subset\Omega_{0}$. Let either $f^{1/(k-1)}\in C^{1,1}(\overline{\Omega_{0}})$ hold or $f^{3/(2k-2)}\in C^{2,1}(\overline{\Omega_{0}})$ and $k\geq 5$ hold. Then, for any constant $\varepsilon>0$, there exists a constant $C_{\varepsilon}>0$ such that
@@
Secondly, the boundary estimate in terms of the interior one for degenerate $k$-Hessian equations is also obtained, but under an additional a priori condition: there exists a small positive constants $r_{0}$ such that
$\Delta u\geq 1\quad\text{in}\ \Omega\cap B_{r_{0}}(y),$ (1.7)
for all $y\in\partial\Omega$ at which $\Delta u(y)\gg 1$. Then the boundary second-order normal derivative estimate is established.
@@
Let $\partial\Omega\in C^{3,1}$ be strictly (k1)(k-1)-convex, $\varphi\in C^{3,1}(\partial\Omega)$, and $f\geq 0$ in $\Omega_{0}$ where $\Omega\Subset\Omega_{0}$. Let one of the following three conditions hold:
[…]
$(iii)\quad f^{3/(2k)}\in C^{2,1}(\overline{\Omega_{0}}),\ \ 2\leq k\leq n-1.$
Then there exists a unique $k$-admissible solution in $C^{1,1}(\overline{\Omega})$ for (1.1).
Remark 1.4 (On condition $(i)$).
If condition (1.7) in $(i)$ could be removed, then the longstanding open problem (1.3) would be resolved. When $\sup_{\partial\Omega}\Delta u$ is bounded from above, the $C^{2}$ estimate follows readily. We hope that Theorem 1.3 may provide some insight into the resolution of the open problem (1.3). Moreover, condition (1.7) is automatically satisfied if $f$ admits a positive lower bound near $\partial\Omega$.
Remark 1.5 (On condition $(iii)$).
As pointed out by Guan-Trudinger-Wang [12], for applications to problems in differential geometry, it is desirable to impose no restriction on the nonnegative function $f$ apart from smoothness. Guan [9] established the existence of a global $C^{1,1}$ solution for the Monge-Ampère equation with homogeneous boundary condition in dimension $n=3$, assuming $0\leq f\in C^{3,1}(\mathbb{R}^{3})$. In comparison with the condition $f^{1/(k-1)}\in C^{1,1}$, it is interesting that the exponent of $f$ can be improved when higher regularity of $f$ is available, for $k\geq 4$.
@@
There is no implication between the conditions $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$ and $f^{1/(k-q)}\in C^{2,1}(\overline{\Omega_{0}})$ for any $0<q<k$. This can be seen from the following two counterexamples. One counterexample is $f^{1/(k-q)}:=|x|^{2}\in C^{\infty}(\overline{\Omega_{0}})$ with $0\in\partial\Omega$. It is clear that $f^{1/k}=|x|^{2(k-q)/k}\notin C^{1,1}(\overline{\Omega_{0}})$. The other counterexample is
@@
in $\overline{\Omega_{0}}$, where 0 and $x_{0}$ are two distinct points on $\partial\Omega$, and
$g(x):=\begin{cases}|x|^{4}\left(1+\sin(1/|x|)\right),\quad x\neq 0,\\ 0,\quad x=0.\end{cases}$
One verifies that $g\in W^{2,\infty}$, hence $g\in C^{1,1}(\overline{\Omega_{0}})$, while $D^{2}g$ is not continuous at 0. Thus $f^{1/k}\in C^{1,1}(\overline{\Omega_{0}})$, but $D^{2}(f^{1/(k-q)})$ fails to be continuous at 0.
The rest of the paper is organized as follows. Section 2 presents preliminary results. Section 3 establishes the global $C^{1}$ estimate and the boundary second-order tangential and tangential-normal estimates, and reduces the estimate of second-order derivatives to the boundary. Section 4 deals with the Monge-Ampère equation under the optimal condition, proving Theorem 1.1. Section 5 is devoted to the weakly interior estimate. Section 6 develops the boundary estimate in terms of the interior one to finish Theorem 1.3.
Throughout the paper, the dependence of constants on the dimension $n$ and the number $k$ in (1.1) is omitted, and a constant depending only on $n$ and $k$ is called universal. This section ends with a list of notations.
@@
10. $D^{2}_{x^{\prime}x^{\prime}}$ denotes a matrix with elements $\partial_{ij}$ where $1\leq i,\ j\leq n-1$.